Oxford Materials IT Committee Meeting

HILARY TERM 2015
A meeting of the Department of Materials IT Committee was held at 11am on Monday 2nd February 2015 in the ETB Committee Room.

AGENDA

  • 1 Apologies for absence and Membership of the Committee
  • 2 Matters outstanding from the previous meeting
  • 3 Matters for discussion (attached)
  • 4 Head of IT report (attached)
  • 5 Finances
  • 6 AOB
  • 7 Date and time of next meeting: Tuesday 12th May 2015 at 11am in PR meeting room

ITEM 2. MATTERS OUTSTANDING FROM THE PREVIOUS MEETING

See minutes below.

ITEM 3: MATTERS FOR DISCUSSION

Verbal reports on the status of various projects:
IT services projects: RDM, SITS/BIZTALK, WDE, TONE, Chorus, ISP
Departmental projects: RexRichards, Photocopiers, iLab
Information from other Committees: MPLS ICT Panel, Divisional Software Licensing Group

ITEM 4: HEAD OF IT REPORT

The report below summarises activities since the last meeting of ITC, as categorised by topic. The report is intended for record rather than for discussion at ITC meeting although queries may arise.

Personnel: Khalid Schofield is still away on unpaid leave since 8th August and there is a possibility he will return at the end of March. Alex continues to provide some support remotely. Numerous tasks have been delayed.

Admin Network Replaced desktop for Jane Foxall. New MacbookAir for Alana Davies and also replaced dual monitors on desktop.

Teaching Laboratory Four of the six brand new teaching lab laptops were replaced after having been flooded by a leak in the heating system (should be reimbursed by estates). Undergraduate printing budgets have been increased to £20 each year in order to cover the increased printing requirements of team design reports and also colour printing of handouts not supplied by lecturers.

Lecture theatres / meeting rooms The HR meeting room AV setup was dismantled and reassembled to allow redecorating of the room. The new chat160 in the Hume-Rothery meeting room seems to work well. Similar Skype functionality is desired for the ETB and BR meeting rooms as part of an upgrade of AV facilities being planned.

The new meeting room in Rex Richards for the Peter Bruce group has taken an extremely long time to specify and order equipment. However installation is imminent on 5th February.

Equalities and Diversity committee have requested electronic display board for visitors both in HR reception and also at Begbroke.

Numerous lectures have been recorded for a student who has been unable to attend lectures due to illness – a new video camera was purchased and the SD cards are regularly sent to the student.

Digital Print-room No significant changes or issues.

Library The self-issue system still needs upgrading. The journal room computer developed an intermittent fan noise and the power supply was replaced.

MML Rothery2 server is still awaiting decommissioning (need to liaise with users of the intel compiler license). Zener is still awaiting reinstallation. Ironman node 30 is no-booting and awaits bios reflashing. Kittel compute nodes were all reinstalled and have been stable for several months.

FIM Several new desktop and reinstallation of older systems. Plans for replacement of image acquisition computer on old 3DAP. Also GDWS to get replacement laptop.

EM EM staff now using laptops. The Begbroke private network was expanded, configurations refined and blocking enabled. SecureShell client software has become incompatible with storage servers prem/bbem (due to using old SSL implementation) therefore WinSCP+putty needs rolling out across the department to replaced SecureShell client.

MFFP Another failed disk replaced in the TJM storage array. Two new storage arrays purchased and all the HPC equipment has been relocated into the HR server room. All systems upgraded to newer versions of Matlab R2014b, Abaqus 6.14-1, Visual Studio 2013 and compatible Intel Fortran Compiler.

OMCS OMCS printer maintenance on oums-bbnano1320n.

Rex Richards / Peter Bruce Group Professor Bruce’s group of 25 people arrived at the start of November. Wireless was implemented ready for their arrival. Phones purchased and operational one week later (on the Cisco VOIP network but ready to move to the Unified VOIP network when it becomes available). PJW registered and configured their computers from St Andrews and supplied new computers on request. By end of January there are 48 PGB group devices on the network. The new Salto door entry system was a mess and incorrectly specified – it still needs fixing to work with university cards rather than separate cards. The meeting room equipment installation is scheduled for 5th February.

Website SSL certificates replaced. New site http://pgbgroup.materials.ox.ac.uk. Content updated for new academic year.

Publications The in-house business card printing service was discontinued and replaced with using Vistaprint to supply business cards. So far 25 people have used the new service – typically £10/100cards. Mugshot display boards were updated for the new academic year and updated again in January. Printed 50 copies of Department Handbook. Weekly editorial of Departmental Newsletter. Signage and posterboards for schools master classes, workshop (SGR) and Hirsch lecture.

Network Network monitoring was reviewed and Nagios, cacti and lantopolog configurations were updated. Rex Richards wireless network was implemented (6 new AP’s) releasing two AP’s to go back into EM IT and HR top floor. Begbroke private network was expanded with some additional cabling. Resolved switch problems with BB Christian, HR stack, ETB core. Tracked down IPv6 broadcast storms affecting the whole network. Improved (but not entirely fixed) networking problems with SAN infrastructure affecting print-server stability.

Security Cotswold security replaced the door access controller for 21BR. Oxcert have released 42 security bulletins since the last meeting including further SSL vulnerabilities. There have been no known compromises. IT staff have dealt with a handful of users with browser hijacks by malware.

WDE Still only six user’ devices encrypted – need to expand uptake targeting senior staff. Chris Akinola discovered a problem with MacOS X Yosemitie upgrade conflicting with WDE.

Disposal WEEE disposal [22 base units; 6 TFT screens; 9 CRT screens; 1 laptop; 20+printers/others Also pneumatically crushed 131 hard disks and 200+ floppy disks] was done for free via a new company called CDL. Another 50+ units are still awaiting disposal.

Software New versions of Matlab/Abaqus/VisualStudio/IntelCompiler was rolled out to all the MFFP workstations.

Disaster Recovery and Backup Numerous disk fails have occurred and data has been restored from backups where possible. A failed disk in the departmental server was replaced under warranty.

Documentation updates Documentation on the private Mediawiki continues to expand, but time for documentation has been limited due to being short staffed.

New Device registrations November 67; December 14; January 9


MINUTES

A meeting of the Department of Materials IT Committee was held at 11am on Monday 2nd February 2015 in the ETB Committee Room. The following people attended: APD, PJW, RS, CA, RT, SCB, JH  

1. Apologies for Absence and Membership of the Committee

Khalid Schofield was out of the country. Jonathan Yates was interviewing.

Lewys Jones was unavailable due to lack of notification about the change of date (sorry).

Flynn Castles has agreed to join the committee but is away this week.

2. Matters arising from the minutes

(Items marked with * were not for discussion at the meeting)

a) PC maintenance – Equinox computer hardware breakdown service (Minute 2a)
A process is still required for (automatic?) notification of owners that their warranty is due to expire. OUCS no-longer offer this service however their supplier Equinox continues to offer cost effective insurance if warranty is required – such as for laptops. No progress.

This item has been on the agenda for several years and is unlikely to be resolved in the near future therefore after some discussion it was agreed to drop this item from the minutes.

b) *Photocopying and Printing (Minute 2b)
All networked printers have been created in the new Print server and the drivers tested. Instructions for installing printers have been updated for the new print server. The old print server is almost ready to turn off. Users will be notified via the Newsletter and via email about a changeover day which may require users to reinstall any printers still on the old iPrint server.

Action: PJW

c) Draft Information Security Policy (Minute 2c)
Again there has been little opportunity to progress towards formalising the departmental Information Security Policy for approval by the ITC and subsequently DC. However PJW noted that the ISP documentation and toolkit has evolved considerably over the last few months and PJW is keen to revisit implementing a departmental IS policy.

Action: PJW

d) XP end of life (Minute 2e)
As of 2nd February 2015 there are still 14 XP systems on the public network.

Action: PJW

e) Office 2003 end of life planning (Minute 2f)
As of 2nd February 2015 there are still 19 computers running Office2003 of which only 6 are on the public network.

Action: PJW

f) *Fresher survey 2013 (Minute 2i)
The six access points (£200/AP) were installed in Rex Richards releasing two existing access points for reuse in EM IT and HR top floor.

3 Matters for Discussion

a) Verbal update on various projects

PJW updated the committee verbally upon the status of the following projects:

  • RDM http://researchdata.ox.ac.uk PJW noted that the ESPRC guidelines on publishing all research data come into force on 15th May 2015. MPLSICT panel were briefed that the university seems likely to define “data” as just the data supporting publications rather than all research data produced by a project. Since it is a university-level obligation to EPSRC (rather than individuals) there is a central push towards compliance and systems are being put in place with training being offered.
  • eVision/SITS/BIZTALK – first phase of the new student records system went live early January without any noticeable problems. http://www.admin.ox.ac.uk/studentsystems/
  • WDE – encryption of laptops is continuing slowly. Note WDE conflicts with MacOS upgrade to Yosemite. http://help.it.ox.ac.uk/wde/index
  • TONE – replacement backbone network to be implemented by end June 2015. Migration of all buildings and replacement of all FroDo units by April 2016. http://projects.it.ox.ac.uk/tone/
  • Chorus – the new telephone system is undergoing acceptance testing and a pilot test will start before Easter. http://projects.it.ox.ac.uk/icp/
  • Innovation Seed Fund – Money is available for small start-up digital projects that will benefit the university. The current round is specifically foccussed on initiatives that will improve the academic experience at Oxford, widen access and improve Student Feedback. Browse and contribute ideas at https://oxfordideas.wazoku.com/. PJW asked JH to raise awareness amongst graduate students.
  • C3E – the university has been reviewing what might be involved if the university were to move to cloud based email solution (e.g. outlook.com / gmail.com) as opposed to replacing Nexus in-house. https://sharepoint.nexus.ox.ac.uk/sites/itservices/c3e

Departmental projects:

  • RexRichards - Professor Bruce’s group of 25 people arrived at the start of November. 48 network registrations so far. Wifi and phones operational. Group website operational. Meeting-room audio visual equipment being installed soon. Salto door locks not as expected and still need changing. This has been a significant time commitment.
  • Photocopiers – Les Chorley and APD and PJW are working towards taking out a new photocopier support contract, getting replacement colour photocopiers with networked print-on-release functionality.
  • iLab – The system has been running for several months. JH reported that iLab booking via mobile phone seems to work now. PJW noted that we are awaiting feedback from iLab about why a few user cost code details appeared to have changed. APD noted that there is still a desire to implement the usage tracker on EM systems. PJW noted that he had rationalized the old equipment booking system at the start of 2015.

4. Head of IT Report
A written report was circulated with the agenda.

5. *Finances
No information was presented at the meeting. Paul still needs to journal costs for stock computers from August onwards.

6. AOB

a) Hotdesk provision for Begbroke users visiting central site
Users from Begbroke visiting central site have asked for access to a hot desk computer. [Users from central site visiting Begbroke are able to use the hotdesk pc provided in the Yarnton Meeting room in Hirsch South Extension.] Current open-access computer provision on central site is:

  • the two library computers in HR (primarily for literature searching – unclear how frequently these are used – all hours access)
  • the two digital printroom computers in ETB room 10.14 (primarily for scanning and printing – often used – open access during office hours plus out of hours access for registered swipecards)
  • the 12 computers in the HB teaching laboratory computer room (primarily for use by undergraduates – access only during lab hours).

After some discussion about access it was agreed that the department provide a new computer in the library as a “hotdesk” computer – running Novell login authentication using departmental username/password and also configured for printing to the Hume-Rothery 5550 printer on the top floor. PJW agreed to notify JH and advertise via newsletter when it is available.

Action: PJW/RS

b) Mac Management
PJW noted that he is actively investigating an improved method for asset management of departmentally owned Apple computers.

c) Apple TV
SCB asked whether it would be possible for the AppleTV+wirelessrouter used to connect to the data projector in the PR meeting room to be given internet access. IT staff to investigate.

Action: PJW / RS / CA

7. Date of Next Meeting
The next meeting is on Tuesday 12th May 2015 at 11am in PR meeting room.

The meeting closed at 12.05pm.